OpenStudy (king2001):

For the function given, state the starting point for a sample period:f(t)=-100sin(50t-20)

RhondaSommer (rhondasommer):

\[\Huge\color{#480a99}{\mathbb{WELCOME\\TO\\OPENSTUDY}}\]

OpenStudy (king2001):

answer choices are (30,0), (2/5,100), (20,0),(2/5,0)

OpenStudy (abdullahm):

f(t) = sin(t) starts at (0, 0) Even with an amplitude of 100, it will start at (0, 0) We now have to consider the phase shift. y = Asin(Bx - C) + D The phase shift is C/B What is the phase shift in your function?

OpenStudy (king2001):

i don't even know what to do. can u answer it for me? i have a test due and i need to pass it

OpenStudy (abdullahm):

What you can easily do is graph it: https://www.desmos.com/calculator/gjek9vojsn Where does it start at?

OpenStudy (king2001):

does it start at 0,0

OpenStudy (abdullahm):

No, it doesn't. That is where it would start at if we didn't have any phase shifts.

OpenStudy (king2001):

i don't know give me hint

RhondaSommer (rhondasommer):

I'm sorry but we're not allowed to help on tests/quizzes/exams/assessments here :/ that would be considered cheating :( http://pages.citebite.com/c3n6a1q0a0mvl

OpenStudy (abdullahm):

@king2001 Looking at the graph, what is the first place it crosses the x-axis? https://www.desmos.com/calculator/gjek9vojsn

OpenStudy (king2001):

the 10?

OpenStudy (abdullahm):

http://prntscr.com/ayhd95

OpenStudy (king2001):

ok whats next then

OpenStudy (abdullahm):

That is the point.

OpenStudy (king2001):

ok

OpenStudy (king2001):

so what would be the answer then

OpenStudy (abdullahm):

Look at your answer choices. All you need to do is convert 0.4 into a fraction.

OpenStudy (king2001):

that would be 2/5

OpenStudy (king2001):

would it be 2/5,100 right or 2/5, 0

OpenStudy (abdullahm):

The point in the graph says (0.4, 0) The y-coordinate is 0. Which one of those two has 0 as the y-coordinate?

OpenStudy (king2001):

2/5,0

OpenStudy (abdullahm):

Correct

OpenStudy (king2001):

thanks. i have one more problem can u help me with it

OpenStudy (abdullahm):

Sure! Make a new post for it!

OpenStudy (king2001):

ok thx